# Embodied Agents Meet Personalization: Investigating Challenges and Solutions Through the Lens of Memory Utilization

### v2

@@ Abstract
LLM-powered embodied agents have shown success on conventional object-rearrangement tasks, but providing personalized assistance that leverages user-specific knowledge from past interactions presents new challenges. We investigate these challenges through the lens of agents’ memory utilization along two critical dimensions: object semantics (identifying objects based on personal meaning) and user patterns (recalling sequences from behavioral routines). To assess these capabilities, we construct Memento, an end-to-end two-stage evaluation framework comprising single-memory and joint-memory tasks. Our experiments reveal that current agents can recall simple object semantics but struggle to apply sequential user patterns to planning. Through in-depth analysis, we identify two critical bottlenecks: information overload and coordination failures when handling multiple memories. Based on these findings, we explore memory architectural approaches to address these challenges. Given our observation that episodic memory provides both personalized knowledge and in-context learning benefits, we design a hierarchical knowledge graph-based user-profile memory module that separately manages personalized knowledge, achieving substantial improvements on both single and joint-memory tasks. Project website: https://connoriginal.github.io/MEMENTO
@@ 2 Related Work
Personalization for embodied agents. The importance of personalization in robotics has long been recognized (Dautenhahn, 2004; Lee et al., 2012; Clabaugh & Matarić, 2018), particularly in the context of human-robot interaction where robots adapt their interactive behaviors to align with individual users. Recent works have focused on reflecting individual user’s preferences during embodied agents’ task execution, such as spatial arrangement (Kapelyukh & Johns, 2021; Wu et al., 2023), table settings (Puig et al., 2021), or personalized object navigations (Dai et al., 2024; Barsellotti et al., 2024). Recently, Xu et al. (2024) aim to infer user preferences from a few demonstrations and adapt planning behavior accordingly. However, these approaches primarily address implicit preference adaptation or short-term reactive behaviors derived from repeated user demonstrations, without leveraging explicitly provided, user-defined knowledge in a structured manner.
@@ 3.1 Task & Focus
Task: personalized object rearrangement. We formulate the personalized object rearrangement task as a Partially Observable Markov Decision Process (POMDP) $(S,A,T,R,\Omega,O,\gamma)$, where $S$ denotes the set of environment states, $A$ is the set of actions, $T$ is the transition function, $R$ is the reward function, $\Omega$ is the observation space, $O$ is the observation function, and $\gamma$ the discount factor. At timestep $t$, the agent receives a text observation $w_{t}\in\Omega$ describing nearby objects and generates actions via LLM policy $\pi$. Given a natural language instruction $I$ that may contain personalized references (e.g., "Place my favorite mug on the table"), the agent must derive the goal representation $g$ using both the instruction and personalized knowledge from memory: $\phi(I,M)\longrightarrow g={(o_{i},l_{i})}_{i=1}^{k}$, where $M={h_{1},h_{2},\ldots,h_{n}}$ represents the set of episodic memories from previous interactions, and each $(o_{i},l_{i})$ pair specifies target object $o_{i}$ and location $l_{i}$. The policy generates actions based on the observation-action trajectory $\tau_{t}$:
@@ 3.4 Metrics & Retrieval Setup
Retrieval setup. For memory retrieval in the utilization stage, we use top-$k$ similarity-based retrieval with all-mpnet-base-v2 (Reimers & Gurevych, 2019), treating the current instruction as a query and episodic memory instructions as keys. Since our primary objective is evaluating memory utilization capability rather than retrieval performance, we ensure the corresponding gold memories are always included in the top-$k$ results through randomly replacing retrieved candidates when necessary. We set $k=5$ as our default configuration because this achieves 96.5% recall for gold memories in realistic retrieval scenarios, compared to only 86.1% recall at $k=3$.
@@ 5.1 Analyzing Agent Performance Under Increased Memory Context
Information overload degrades memory utilization performance. As shown in Figure 6, increasing $k$ consistently degrades performance across all models and both personalized knowledge types, indicating that agents struggle to extract relevant information when presented with larger sets of retrieved memories. We found that agents increasingly rely on commonsense knowledge rather than personalized memories as retrieved memories increase, especially on user patterns, suggesting that embodied agents suffer from information overload. Therefore, even though LLMs demonstrate impressive capabilities in processing long-context information (Wang et al., 2024b; Liu et al., 2025), memory systems that retrieve appropriate numbers of memories are crucial for personalized embodied agents. We provide further analysis of retrieval strategies in scenarios without guaranteed gold memory access in Appendix E.3.
@@ 7 Conclusion
This work involves no human subjects, conflicts of interest, or legal compliance concerns. We use licensed resources (e.g., PARTNR (Chang et al., 2024) under the MIT License) and commonsense knowledge without personal data, copyrighted content, or bias-amplifying material. We generated personalized knowledge from LLM-synthesized descriptions with prompts designed to reduce biased or discriminatory content, and we manually verified that the resulting user knowledge was free of such concerns. We acknowledge that embodied memory systems may pose privacy and security risks in real-world deployment. To mitigate these issues, future systems should adopt on-device processing, minimize retention, require user consent, and apply secure, transparent memory handling. We will release only de-identified or synthetic assets and provide documentation of safeguards to reduce misuse risk.
@@
To facilitate full reproducibility of our results, we include complete evaluation dataset details and experimental procedures in the appendices. Furthermore, we publish our code at https://anonymous.4open.science/r/MEMENTO with thorough documentation.
@@ A.3 Societal Impacts
Our work explores how embodied agents can remember and adapt to user-specific preferences through episodic memory, enabling more personalized and natural interactions. This capability has the potential to enhance convenience, efficiency, and user satisfaction in everyday environments—benefiting a wide range of users regardless of age or ability. However, personalization also raises concerns about privacy, bias reinforcement, and over-dependence on AI agents (Rueben & Smart, 2016; Lee et al., 2012). Since episodic memory involves storing interaction history, future systems must consider secure and transparent memory handling. Although our study is conducted in a controlled simulator, these considerations will become crucial as such systems move toward real-world deployment. We hope that Memento serves as a foundation for future research in building safe, privacy-aware, and user-aligned personalized embodied agents.
@@ C.1 LLM-Powered Embodied Agent Architecture
For retrieval, we encode instructions and memory entries using the all-mpnet-base-v2 sentence transformer (Reimers & Gurevych, 2019) and use the current task instruction as the query. To avoid ambiguous or irrelevant memories, we retrieve candidate memories only from the history within the same scene as the current task.
@@ D.1 User Profile Memory
// Step 5: Use LLM to decide add or update
$op\leftarrow LLMResolve(I,C)$;
@@
Knowledge graph update procedure. While the hierarchical design specifies how knowledge is organized, the graph must also evolve as users provide new information or refine existing entries. Algorithm 1 presents the procedure for updating the knowledge graph. Given a user instruction $I$ and the current graph $G$, the procedure begins by extracting candidate knowledge $\mathcal{K}$ and associated elements $\mathcal{E}$. For each candidate knowledge $k$, the relevant subgraph $G_{t_{k}}$ is filtered, and similarity search is applied to collect matching nodes $C$. These candidates are then expanded by collecting all connected edges and their descendant nodes, and the resulting subgraphs are passed to an LLM. The LLM determines whether the instruction corresponds to an addition or an update and performs the corresponding operation, yielding the updated graph $G$’.
@@
Knowledge graph schema. We summarize the node and edge types in Tables 7 and 8. Complete examples of user profile memory instances are shown in Figures 12 and 13.
@@ Appendix E Extended Analysis: Memory Components and Real-World Scenarios
E.1 Human Baseline
Experiment setup.
  • Easy: Llama-8B successfully completed the task using its own memory
  • Hard: Llama-8B failed even when given GPT-4o’s answer memory
  • Super hard: GPT-4o failed with its own memory
  • Random joint-memory tasks
Results.
E.2 Memory Quality and Robustness Analysis
To better understand the factors affecting memory utilization performance, we conduct additional analysis examining two critical aspects of memory-based personalized assistance. First, we analyze how memory quality impacts agent performance by comparing gold memory trajectories with retrieved memory from actual agent interactions. Second, we evaluate agent robustness when episodic memories are corrupted or incomplete through various noise and degradation conditions. These analyses provide insights into the practical limitations and requirements for deploying personalized embodied agents in real-world scenarios where memory quality cannot be guaranteed.
Figure 11: The results of memory quality analysis.
E.2.1 Memory Quality Analysis
Experiment setup.
We further analyze the effect of memory quality by comparing gold memory, consisting of successful and shortest-path trajectories that serves as high-quality references with the memory obtained from interaction histories in the declaration stage. By evaluating performance differences between these two settings, we aim to understand how memory quality affects agent performance across the memory utilization stage for both tasks (single-memory tasks, joint-memory tasks).
Performance degradation with lower-quality trajectories.
Figure 11 shows the performance comparison between gold memory and retrieved memory across the memory utilization stage. In the memory utilization stage, high-capacity models (Llama-3.1-70b) show relatively stable performance across gold and retrieved memory, while lower-capacity models (Llama-3.1-8b) exhibit a substantial drop when using retrieved memory. This suggests that less capable models have more difficulty extracting relevant information from an imperfect memory context. When executing more demanding joint-memory tasks, which require combining and reasoning across multiple memory sources, performance degrades sharply across all models with retrieved memory. This indicates that the challenge of integrating multiple memories introduces compounding complexity, as the agent must not only retrieve relevant information but also correctly synthesize disparate memory contexts. These findings emphasize a fundamental limitation of retrieval-based memory systems; while gold memory provides ground truth references, retrieved memory is inherently prone to semantic noise, as it relies on approximate similarity rather than precise matching. Enhancing memory quality through more precise filtering is essential to improve reasoning performance, particularly for complex joint tasks that require integration of multiple memory sources, and especially for smaller models with limited reasoning capabilities.
E.2.2 Robustness to Noisy and Partial Memories
[…]
GPT-4o Gold Baseline 14.5 2481.2 85.3 80.6
Noisy Random 14.6 2759.0 81.5 -3.8 78.6 -2.0
Shuffle 14.4 2450.9 84.0 -1.3 80.6 0.0
Partial Random 14.5 2526.9 84.4 -0.9 80.1 -0.5
w/o Pick 14.6 2555.7 82.9 -2.4 79.1 -1.5
w/o Place 14.5 2569.8 84.4 -0.9 82.6 +2.0
w/o Nav 15.4 2526.8 81.4 -3.9 78.5 -2.1
Experiment setup.
Random noise and missing navigation cues degrade performance.
E.3 Retrieval Strategy Analysis
Table 10: We use an embedding-based retrieval method with the all-mpnet-base-v2 Sentence Transformer (Reimers & Gurevych, 2019).
Retriever Strategy Top-K Missing Rates (%)
Baseline 3 13.9
5 3.5
E.3.1 Baseline Retrieval Analysis.
To assess whether the corresponding episodic memory is successfully retrieved during memory utilization, we first measure the missing rate under a baseline similarity-based retriever.  [^1] We adopt an embedding-based method using the all-mpnet-base-v2 Sentence Transformer (Reimers & Gurevych, 2019). The results show missing rates of 13.9% for top-3 and 3.5% for top-5, indicating that expanding to top-5 nearly eliminates missing cases. Based on this observation, we set top-5 as the default retrieval setup, prioritizing inclusion of the required memory over the risk of distractor entries.
E.3.2 Advanced Retrieval Strategy Analysis
GPT-4o Baseline 3 13.9 82.9 79.5
5 3.5 84.7 82.1
Reranker 3 2.5 87.2 84.1
5 1.5 84.3 80.6
Query Expansion 3 9.0 84.9 81.1
5 1.5 86.2 81.1
Qwen-2.5-72b Baseline 3 13.9 75.3 70.1
5 3.5 75.4 70.0
Reranker 3 7.6 73.0 67.2
5 2.1 70.9 64.1
Query Expansion 3 8.5 75.6 69.2
5 4.0 72.6 66.2
Experiment setup.
Improved retrieval does not ensure improved outcomes.
E.4 Real-World Scenario Analysis
E.4.1 Real-World Multi-Episode Evaluation
Table 12: Performance results on multiple joint-memory tasks.
Model Task PC (%) SR (%)
GPT-4o joint memory task 86.7 63.9
GPT-4o 3-joint memory task 61.2 22.0
Qwen-2.5-72b joint memory task 68.9 36.1
Qwen-2.5-72b 3-joint memory task 42.4 7.3
E.4.2 Agent Behavior under Ambiguous Instructions from Users
Model PC (%) SR (%)
GPT-4o (Baseline) 92.0 90.0
GPT-4o 80.4 73.3
Qwen-2.5-72b (Baseline) 75.1 66.7
Qwen-2.5-72b 59.6 53.3
Experiment setup.
Result.
E.4.3 Personalized Knowledge Generalization
Experiment setup.
Model Task PC (%) SR (%)
GPT-4o Baseline 100.0 100.0
GPT-4o Generalization 92.5 83.3
Qwen-2.5-72b Baseline 75.0 75.0
Qwen-2.5-72b Generalization 85.8 75.0
Result.
Appendix F Prompts
F.1 Prompts for Dataset Generation
F.2 Prompts for Agent
F.3 Prompts for Memory Design Experiment

### v3

@@ Abstract
LLM-powered embodied agents have shown success on conventional object-rearrangement tasks, but providing personalized assistance that leverages user-specific knowledge from past interactions presents new challenges. We investigate these challenges through the lens of agents’ memory utilization along two critical dimensions: object semantics (identifying objects based on personal meaning) and user patterns (recalling sequences from behavioral routines). To assess these capabilities, we construct Memento, an end-to-end two-stage evaluation framework comprising single-memory and joint-memory tasks. Our experiments reveal that current agents can recall simple object semantics but struggle to apply sequential user patterns to planning. Through in-depth analysis, we identify two critical bottlenecks: information overload and coordination failures when handling multiple memories. Based on these findings, we explore memory architectural approaches to address these challenges. Given our observation that episodic memory provides both personalized knowledge and in-context learning benefits, we design a hierarchical knowledge graph-based user-profile memory module that separately manages personalized knowledge, achieving substantial improvements on both single and joint-memory tasks. Our code and data is available at https://github.com/Connoriginal/MEMENTO.
@@ 2 Related Work
Personalization for embodied agents. The importance of personalization in robotics has long been recognized (Dautenhahn, 2004; Lee et al., 2012; Clabaugh and Matarić, 2018), particularly in the context of human-robot interaction where robots adapt their interactive behaviors to align with individual users. Recent works have focused on reflecting individual user’s preferences during embodied agents’ task execution, such as spatial arrangement (Kapelyukh and Johns, 2021; Wu et al., 2023), table settings (Puig et al., 2021), or personalized object navigations (Dai et al., 2024; Barsellotti et al., 2024). Recently, Xu et al. (2024) aim to infer user preferences from a few demonstrations and adapt planning behavior accordingly. However, these approaches primarily address implicit preference adaptation or short-term reactive behaviors derived from repeated user demonstrations, without leveraging explicitly provided, user-defined knowledge in a structured manner.
@@ 3.1 Task & Focus
Task: personalized object rearrangement. We formulate the personalized object rearrangement task as a Partially Observable Markov Decision Process (POMDP) $(S,A,T,R,\Omega,O,\gamma)$, where $S$ denotes the set of environment states, $A$ is the set of actions, $T$ is the transition function, $R$ is the reward function, $\Omega$ is the observation space, $O$ is the observation function, and $\gamma$ the discount factor. At timestep $t$, the agent receives a text observation $w_{t}\in\Omega$ describing nearby objects based on the previous action $a_{t-1}$ and generates action $a_{t}$ via LLM policy $\pi$. Given a natural language instruction $I$ that may contain personalized references (e.g., "Place my favorite mug on the table"), the agent must derive the goal representation $g$ using both the instruction and personalized knowledge from memory: $\phi(I,M)\longrightarrow g={(o_{i},l_{i})}_{i=1}^{k}$, where $M={h_{1},h_{2},\ldots,h_{n}}$ represents the set of episodic memories from previous interactions, and each $(o_{i},l_{i})$ pair specifies target object $o_{i}$ and location $l_{i}$. The policy generates actions based on the observation-action trajectory $\tau_{t}$:
@@ 3.4 Metrics & Retrieval Setup
Retrieval setup. We first run the memory acquisition stage for all episodes to accumulate episodic memories, which are then used in the memory utilization stage. For memory retrieval in the utilization stage, we use top-$k$ similarity-based retrieval with all-mpnet-base-v2 (Reimers and Gurevych, 2019), treating the current instruction as a query and episodic memory instructions as keys. Since our primary objective is evaluating memory utilization capability rather than retrieval performance, we ensure the corresponding gold memories are always included in the top-$k$ results through randomly replacing retrieved candidates when necessary. We set $k=5$ as our default configuration because this achieves 96.5% recall for gold memories in realistic retrieval scenarios, compared to only 86.1% recall at $k=3$. Comprehensive analysis of various retrieval strategies without gold memory guarantees is provided in Appendix E.3.
@@ 5.1 Analyzing Agent Performance Under Increased Memory Context
Information overload degrades memory utilization performance. As shown in Figure 6, increasing $k$ consistently degrades performance across all models and both personalized knowledge types, indicating that agents struggle to extract relevant information when presented with larger sets of retrieved memories. We found that agents increasingly rely on commonsense knowledge rather than personalized memories as retrieved memories increase, especially on user patterns, suggesting that embodied agents suffer from information overload. Therefore, even though LLMs demonstrate impressive capabilities in processing long-context information (Wang et al., 2024b; Liu et al., 2025), which we confirm in Appendix D.5, memory systems that retrieve appropriate numbers of memories are crucial for personalized embodied agents.
@@ 7 Conclusion
Acknowledgements
This project is supported by Microsoft Research Asia. This research was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT), under the Global Research Support Program in the Digital Field program (RS-2024-00436680), National AI Research Lab Project (RS-2024-00457882), Artificial Intelligence Graduate School Program (Yonsei University) (RS-2020-II201361). We also thank Donghyun Kim, Juhyun Park, Sungyong Park, and Hyeonjong Ju for their helpful discussions and support. Jinyoung Yeo is the corresponding author.
@@ Acknowledgements
This work involves no human subjects, conflicts of interest, or legal compliance concerns. We use licensed resources (e.g., PARTNR (Chang et al., 2024) under the MIT License) and commonsense knowledge without personal data, copyrighted content, or bias-amplifying material. We generated personalized knowledge from LLM-synthesized descriptions with prompts designed to reduce biased or discriminatory content, and we manually verified that the resulting user knowledge was free of such concerns. We acknowledge that embodied memory systems may pose privacy and security risks in real-world deployment. To mitigate these issues, future systems should adopt on-device processing, minimize retention, require user consent, and apply secure, transparent memory handling.
@@
To facilitate full reproducibility of our results, we include complete evaluation dataset details and experimental procedures in the appendices. Furthermore, we publish our code at  https://github.com/Connoriginal/MEMENTO with thorough documentation.
@@ A.3 Societal Impacts
Our work explores how embodied agents can remember and adapt to user-specific preferences through episodic memory, enabling more personalized and natural interactions. This capability has the potential to enhance convenience, efficiency, and user satisfaction in everyday environments—benefiting a wide range of users regardless of age or ability. However, personalization also raises concerns about privacy, bias reinforcement, and over-dependence on AI agents (Rueben and Smart, 2016; Lee et al., 2012). Since episodic memory involves storing interaction history, future systems must consider secure and transparent memory handling. Although our study is conducted in a controlled simulator, these considerations will become crucial as such systems move toward real-world deployment. We hope that Memento serves as a foundation for future research in building safe, privacy-aware, and user-aligned personalized embodied agents.
@@ C.1 LLM-Powered Embodied Agent Architecture
For retrieval, we encode instructions and memory entries using the all-mpnet-base-v2 sentence transformer (Reimers and Gurevych, 2019) and use the current task instruction as the query. To avoid ambiguous or irrelevant memories, we retrieve candidate memories only from the history within the same scene as the current task.
@@ C.3.3 Inter-Rater Reliability Analysis
Figure 11: Distribution of required (object, location) pairs across object semantics and user pattern tasks.
@@ C.3.4 Dataset Statistics
Number of (target, location) pairs. We compared the number of (object, location) pairs required for successful completion across both object semantics and user pattern tasks. As shown in Figure 11, the distribution of required (object, location) pairs is nearly identical between the two personalized knowledge types. This indicates that the performance difference between object semantics and user patterns stems from the intrinsic characteristics of each knowledge type rather than differences in task complexity measured by the number of required object-location pairs.
@@ D.1 User Profile Memory
Knowledge graph update procedure. While the hierarchical design specifies how knowledge is organized, the graph must also evolve as users provide new information or refine existing entries. Algorithm 1 presents the procedure for updating the knowledge graph. Given a user instruction $I$ and the current graph $G$, the LLM processes the instruction by extracting candidate knowledge $\mathcal{K}$ and associated elements $\mathcal{E}$. For each candidate knowledge $k$, the relevant subgraph $G_{t_{k}}$ is filtered, and similarity search is applied to collect matching nodes $C$. These candidates are then expanded by collecting all connected edges and their descendant nodes, and the resulting subgraphs are passed to an LLM. The LLM determines whether the instruction corresponds to an addition or an update, executes the corresponding operation. For an update, the LLM removes the existing knowledge node and generates a new one, and for an addition it simply creates a new knowledge node. In both cases, the LLM determines whether to reuse or instantiate the associated object, location nodes based on whether their similarity scores exceed a predefined threshold.
@@
// Step 5: Decide add or update
$op\leftarrow DecideAddOrUpdate(I,C)$;
@@
Knowledge graph schema. We summarize the node and edge types in Tables 7 and 8 and provide complete examples of user profile memory instances in Figures 12 and 13.
@@ D.4 Example Trajectory and Evaluation
D.5 LLMs Long Context Comprehension Analysis
To investigate whether the information overload bottleneck in LLM-powered embodied agents for personalized assistance stems from limitations in processing long-context information, we conduct an experiment in a QA format to evaluate whether LLMs can properly reference personalized information from long contexts.
Experiment setup. We provide LLMs with the episodic memory information required for personalized assistance tasks along with the corresponding task instructions. We then prompted the models: "Identify the personalized knowledge necessary to perform this instruction based on the context provided below." To evaluate the responses, we use GPT-4o as an LLM-as-a-Judge, supplying it with the stored gold memory and task instruction as references to assess whether the LLM’s identified personalized knowledge aligns with that in the gold memory.
Table 15: Accuracy on long-context QA with varying numbers of retrieved episodic memories (Top-$k$).
Model k = 1 3 5 7 10 20
GPT-4o 100% 98.71% 98.28% 96.98% 96.98% 95.69%
Qwen-2.5-72b 100% 93.97% 94.40% 93.53% 93.10% 83.19%
Results. As shown in the Table 15, all models demonstrate high accuracy in correctly identifying personalized knowledge in the QA format. Compared to the results in Section 5.1 Figure 6, this indicates that current LLMs have the ability to reference personalized information from long contexts, but encounter difficulties when applying this personalized information to planning tasks.
E.1 Human Baseline
Experiment setup.
  • Easy: Llama-8B successfully completed the task using its own memory
  • Hard: Llama-8B failed even when given GPT-4o’s answer memory
  • Super hard: GPT-4o failed with its own memory
  • Random joint-memory tasks
Results.
E.2 Memory Quality and Robustness Analysis
To better understand the factors affecting memory utilization performance, we conduct additional analysis examining two critical aspects of memory-based personalized assistance. First, we analyze how memory quality impacts agent performance by comparing high quality memory trajectories with acquired memory trajectories from actual agent interactions. Second, we evaluate agent robustness when episodic memories are corrupted or incomplete through various noise and degradation conditions. These analyses provide insights into the practical limitations and requirements for deploying personalized embodied agents in real-world scenarios where memory quality cannot be guaranteed.
Table 16: Performance comparison of models when using high-quality memory versus acquired (self-acquired) memory, evaluated in both single- and joint-memory settings.
Model Memory Stage Replanning Count $\downarrow$ Sim Steps $\downarrow$ PC (%) $\uparrow$ SR (%) $\uparrow$
GPT-4o High-quality Single Memory 14.5 2481.2 85.3 80.6
Joint Memory 25.8 3619.4 83.3 63.9
Baseline Single Memory 16.1 2450.8 87.9 85.1
Joint Memory 28.9 3480.7 86.7 63.9
Claude-3.5-Sonnet High-quality Single Memory 14.1 2266.4 77.2 70.1
Joint Memory 24.2 3226.6 69.9 38.9
Baseline Single Memory 15.3 2258.8 69.3 63.7
Joint Memory 27.8 3198.8 64.6 33.3
Qwen-2.5-72B High-quality Single Memory 15.2 2630.4 74.7 68.2
Joint Memory 26.4 4082.1 64.8 38.9
Baseline Single Memory 17.5 2691.2 72.6 67.2
Joint Memory 31.3 4027.1 68.9 36.1
Llama-70B High-quality Single Memory 15.1 2364.0 72.6 67.2
Joint Memory 27.1 3374.5 64.1 27.8
Baseline Single Memory 19.0 2566.6 72.2 66.7
Joint Memory 31.4 3425.2 51.3 8.3
Llama-8B High-quality Single Memory 16.9 2739.0 65.1 59.2
Joint Memory 26.4 3753.2 54.0 13.9
Baseline Single Memory 19.0 3131.7 48.1 35.0
Joint Memory 27.4 3478.2 35.3 8.3
Qwen-2.5-7B High-quality Single Memory 15.7 2991.5 55.7 46.3
Joint Memory 23.9 4402.0 34.0 8.3
Baseline Single Memory 21.8 3271.4 39.1 27.4
Joint Memory 26.9 4148.6 33.7 5.6
E.2.1 Memory Quality Analysis
Experiment setup.
We further analyze the effect of memory quality by comparing high quality memory, consisting of successful and shortest-path trajectories that serves as curated references with the memory obtained from interaction histories in the memory acquisition stage, with predicted memory obtained via automated retrieval. By evaluating performance differences between these two settings, we aim to understand how memory quality affects agent performance across the memory utilization stage for both tasks (single-memory tasks, joint-memory tasks).
Performance degradation with lower-quality trajectories.
Table 16 shows the performance comparison between high quality memory and acquired memory across the memory utilization stage. In the memory utilization stage, high-capacity models show relatively stable performance across gold and retrieved memory, while lower-capacity models exhibit a substantial drop when using retrieved memory. This suggests that less capable models have more difficulty extracting relevant information from an imperfect memory context. When executing more demanding joint-memory tasks, which require combining and reasoning across multiple memory sources, performance degrades sharply across all models with acquired memory. This indicates that the compounded complexity of integrating multiple memories makes it increasingly important for the agent to rely on clear, noise-free, high-quality memory contexts from which synthesis can be performed reliably. Additionally, as shown in the table X, the overall reduction in planning steps when using high-quality memory suggests that episodic memory plays an important role in the planning process of embodied agents. Enhancing memory quality through more precise filtering is therefore essential not only for improving reasoning performance but also for increasing the agent’s planning efficiency.
E.2.2 Robustness to Noisy and Partial Memories
[…]
GPT-4o Gold Baseline 14.5 2481.2 85.3 80.6
Noisy Random 14.6 2759.0 81.5 -3.8 78.6 -2.0
Shuffle 14.4 2450.9 84.0 -1.3 80.6 0.0
Partial Random 14.5 2526.9 84.4 -0.9 80.1 -0.5
w/o Pick 14.6 2555.7 82.9 -2.4 79.1 -1.5
w/o Place 14.5 2569.8 84.4 -0.9 82.6 +2.0
w/o Nav 15.4 2526.8 81.4 -3.9 78.5 -2.1
Experiment setup.
Random noise and missing navigation cues degrade performance.
E.3 Retrieval Strategy Analysis
Table 18: We use an embedding-based retrieval method with the all-mpnet-base-v2 Sentence Transformer (Reimers and Gurevych, 2019).
Retriever Strategy Top-K Missing Rates (%)
Baseline 3 13.9
5 3.5
E.3.1 Baseline Retrieval Analysis.
To assess whether the corresponding episodic memory is successfully retrieved during memory utilization, we first measure the missing rate under a baseline similarity-based retriever.  [^1] We adopt an embedding-based method using the all-mpnet-base-v2 Sentence Transformer (Reimers and Gurevych, 2019). The results show missing rates of 13.9% for top-3 and 3.5% for top-5, indicating that expanding to top-5 nearly eliminates missing cases. Based on this observation, we set top-5 as the default retrieval setup, prioritizing inclusion of the required memory over the risk of distractor entries.
E.3.2 Advanced Retrieval Strategy Analysis
GPT-4o Baseline 3 13.9 82.9 79.5
5 3.5 84.7 82.1
Reranker 3 2.5 87.2 84.1
5 1.5 84.3 80.6
Query Expansion 3 9.0 84.9 81.1
5 1.5 86.2 81.1
Qwen-2.5-72b Baseline 3 13.9 75.3 70.1
5 3.5 75.4 70.0
Reranker 3 7.6 73.0 67.2
5 2.1 70.9 64.1
Query Expansion 3 8.5 75.6 69.2
5 4.0 72.6 66.2
Experiment setup.
Improved retrieval does not ensure improved outcomes.
E.4 Real-World Scenario Analysis
E.4.1 Real-World Multi-Episode Evaluation
Table 20: Performance results on multiple joint-memory tasks.
Model Task PC (%) SR (%)
GPT-4o joint memory task 86.7 63.9
GPT-4o 3-joint memory task 61.2 22.0
Qwen-2.5-72b joint memory task 68.9 36.1
Qwen-2.5-72b 3-joint memory task 42.4 7.3
E.4.2 Agent Behavior under Ambiguous Instructions from Users
Model PC (%) SR (%)
GPT-4o (Baseline) 92.0 90.0
GPT-4o 80.4 73.3
Qwen-2.5-72b (Baseline) 75.1 66.7
Qwen-2.5-72b 59.6 53.3
Experiment setup.
Result.
E.4.3 Personalized Knowledge Generalization
Experiment setup.
Model Task PC (%) SR (%)
GPT-4o Baseline 100.0 100.0
GPT-4o Generalization 92.5 83.3
Qwen-2.5-72b Baseline 75.0 75.0
Qwen-2.5-72b Generalization 85.8 75.0
Result.
E.4.4 User Profile Memory Robustness
Experiment setup.
In real-world interactions, users often refer to knowledge, objects, or locations using different names or expressions rather than clean and consistent symbolic identifiers. To directly evaluate the robustness of our proposed User Profile Memory under realistic conditions, we additionally conduct a qualitative analysis on how well User Profile Memory handles natural noise that arises in real user interactions. We identified two noise types that naturally arise when users interact with a personalized embodied agent:
  • Knowledge Update Noise - the user expresses an existing knowledge using a different phrase (synonyms, paraphrases, alias names).
  • Node Reference Noise - new knowledge refers to an existing object / location using a different form.
For each category, three undergraduate annotators modified the original instructions grounded in the scene of Memento, producing 10 noisy instructions for each noise type. This procedure ensures the noise realistically reflects how humans describe object semantics or user patterns. These noisy instructions were crafted using the underlying scene (103997895_171031182103997895\_171031182), and to ensure an accurate evaluation of noise robustness, all experiments were conducted on graphs constructed from the original instructions. Full datasets are provided in Table 24 and Table 25
Table 24: Dataset of Knowledge Update Noise Variants.
Original Knowledge Modified Knowledge
Gift teapot from grandfather Grandfather’s gifted teakettle
Family heirloom vase Inherited family vase
Gift toy airplane from childhood friend A toy airplane that a childhood friend gave me
Favorite relaxation candle The candle I always use to relax
Collectible android figure Android collectible figurine
Bedroom organization Organizing setup for the bedroom space
Morning work session setup Morning workspace arrangement
Cozy room setup Arrangement for creating a cozy room atmosphere
Bedtime setup Nighttime setup before going to bed
Study materials organization Organized placement of study materials
Table 25: Dataset of Node Reference Noise Variants.
Original Description Modified Description
Black teapot with curved spout and ornate lid A curved-spout black teapot featuring an intricately decorated lid
Gray laptop with black keyboard/touchpad A gray laptop equipped with a black keyboard and touchpad
Dark green analog alarm clock with twin bells Dark green twin-bell analog alarm clock
Black and orange keychain with circular pendant A circular pendant keychain in black and orange
android figure which is a black and white android panda figure a black and white Android panda figurine
Move the toy airplane from the bedroom to the living room table. Place it next to the couch. Relocate the toy plane from the bedroom onto the table in the living room, positioning it next to the sofa.
Move the picture frame and clock to the table in the bedroom. Then, place them next to each other on the counter in the kitchen. Bring the photo frame and the alarm clock to the bedroom table, and afterward arrange both items side by side on the countertop at kitchen.
Place the bread, the potato, and the egg on the kitchen table and place them next to each other. Set the loaf, spud, and fresh egg together on the table at kitchen, arranging them close to one another.
Move the laptop and the mouse pad to a new living room table. Move the clock to another shelf. Transfer the notebook computer and mouse mat onto the new table in the living room, and place the clock on a separate shelf.
Take the book, pencil case, and multiport hub from the table in the living room and put them on the shelves in the bedroom, placing them next to each other. Gather the note, pencil pouch, and usb hub from the living room table and arrange them side by side on the bedroom shelves.
Result.
Table 26: Accuracy under different noise types.
Noise Type Accuracy (%)
Knowledge Update Noise 100.0
Node Reference Noise 75.0
[…]
Table 27: Failure cases in node reference noise
Existing Node Add Node
mouse pad mouse mat
clock alarm clock
android figure which is a black and white android panda figure black and white Android Panda Figurine
kitchen counter countertop at kitchen
another shelf separate shelf
Appendix F Prompts
F.1 Prompts for Dataset Generation
F.2 Prompts for Agent
F.3 Prompts for Memory Design Experiment